# Enhancing Perception Quality in Remote Sensing Image Compression via Invertible Neural Network

### v1

@@ Abstract
Decoding remote sensing images to achieve high perceptual quality, particularly at low bitrates, remains a significant challenge. To address this problem, we propose the invertible neural network-based remote sensing image compression (INN-RSIC) method. Specifically, we capture compression distortion from an existing image compression algorithm and encode it as a set of Gaussian-distributed latent variables via INN. This ensures that the compression distortion in the decoded image becomes independent of the ground truth. Therefore, by leveraging the inverse mapping of INN, we can input the decoded image along with a set of randomly resampled Gaussian distributed variables into the inverse network, effectively generating enhanced images with better perception quality. To effectively learn compression distortion, channel expansion, Haar transformation, and invertible blocks are employed to construct the INN. Additionally, we introduce a quantization module (QM) to mitigate the impact of format conversion, thus enhancing the framework’s generalization and improving the perceptual quality of enhanced images. Extensive experiments demonstrate that our INN-RSIC significantly outperforms the existing state-of-the-art traditional and deep learning-based image compression methods in terms of perception quality.
1 Introduction
@@
Remote sensing (RS) images are widely used in many fields, including urban planning [1], biosphere monitoring [2], and land resource management [3]. However, RS images are typically characterized by high content, high resolution, and large size [4, 5]. Furthermore, with the advancement of sensor technology and the enhancement of the image acquisition capability of satellite and airborne equipment, the necessity of storing or transmitting RS images is also growing in importance. To cope with these challenges, high-resolution RS images are usually compressed before they are stored and transmitted to the ground. In this case, it is critical to achieve low-bitrate compression while obtaining the decoded images with high perceptual quality.
@@
Traditional image compression standards, such as JPEG2000 [6], BPG [7], and VVC [8], have been instrumental in facilitating the storage and transmission of image data. However, image compression with these algorithms inevitably suffered from undesired blocking, ringing artifacts, and blurring [9, 10], which may highly influence the perception quality.
@@
With the development of deep learning techniques, learning-based image compression methods have made significant progress in obtaining high rate-distortion performance [11, 12, 9, 13]. Notably, in natural image compression, Ballé et al. [14] introduced additional side information as a hyper-prior entropy model to estimate a zero-mean Gaussian distribution, laying the foundation for subsequent improvements in entropy modeling. Cheng et al. [10] further proposed the use of discretized Gaussian mixture likelihoods to parameterize the distributions of the entropy model, resulting in impressive decoding performance. Meanwhile, He et al. [15] proposed an innovative uneven channel-conditional adaptive grouping scheme, enhancing compression performance without sacrificing computational speed. Despite the success of learning-based compression algorithms in natural scenes, RS images present unique challenges due to their rich texture, weak correlation, and low redundancy compared to natural images [16, 17]. Preserving visually pleasing decoded images while achieving high compression bitrates remains a significant challenge in RS image compression.
@@
To tackle the challenge, Zhang et al. [11] introduced a multi-scale attention module to enhance the network’s feature extraction capability, and leveraged global priors and anchored-stripe attention to improve the entropy model. Pan et al. [18] employed generative adversarial networks (GANs) to independently decode the image content and detailed textures, subsequently fusing these features to achieve low-bitrate RS image compression. Additionally, Xiang et al. [19] utilized discrete wavelet transform to separate image features into high- and low-frequency components, and designed compression networks to enhance the model’s representation ability of both types of features.
@@
In summary, the aforementioned deep learning-based image compression methods typically achieve image compression through CNN blocks [10, 15, 14, 13, 11, 19] or Transformer blocks [20, 21, 22, 23, 24], followed by optimization using distance measurements like L1, L2, or adversarial loss. Despite their impressive performance, these methods usually struggle to generate satisfactory perception quality. Recently, invertible neural network (INN) has emerged as a new paradigm for image generation, demonstrating remarkable performance across various applications [25, 26, 27, 28]. Building on this success, several approaches have integrated INN into image compression to capture richer texture details [23, 26]. For instance, in [23], an enhanced INN-based encoding network was devised to improve the decoding performance of natural image compression. In [26], the authors utilized INN to develop an invertible image generation module, aiming to prevent information loss and propose a competitive low-bit-rate compression algorithm. In contrast to these methods, our approach focuses on leveraging the invertible capabilities of INN to recover more texture information of images decoded by an existing compression algorithm, without requiring additional bitstreams and model retraining. It is expected to sample some texture details from Gaussian distributed variables conditioned on the decoded image. Figure 1 illustrates the perception comparison between the decoded images and the enhanced images. It is observed that our INN-RSIC exhibits a high capability in recovering texture-rich details without additional bitrates.
Figure 1: Visualization of the decoded results by ELIC [15] and our INN-RSIC.
@@
In this paper, we propose an INN-based RS image compression (INN-RSIC) method. Specifically, we utilize an invertible forward network with conditional generation module (CGM) to encode compression loss information of an existing image compression algorithm into Gaussian-distributed latent variables. As a result, it is expected to sample some of the lost prior information from Gaussian distribution, thereby facilitating the recovery of visually enhanced images during the inverse mapping. Additionally, to effectively learn compression distortion, we adopt channel expansion and Haar transformation [29] to separate high and low frequency and introduce a quantization module (QM) to reduce the impact of the reconstruction quality due to data type conversion in the inference stage. The primary contributions of this paper can be summarized as follows: 1) To the best of our knowledge, our INN-RSIC is the first attempt to model image compression distortion information of an existing image compression algorithm using invertible transforms. 2) We develop a novel, effective yet simple architecture with channel expansion, Haar transform, and invertible blocks. This architecture enables projecting compression distortion into a case-agnostic distribution, so that compression distortion information can be obtained based on samples from a Gaussian distribution. 3) CGM is introduced to split and encode the compression distortion information from ground truth into the latent variables conditioned on the synthetic image. This process generates pattern-free synthetic images and enhances the texture details of the reconstructed images during the inverse mapping. 4) Extensive experiments indicate that our INN-RSIC outperforms the existing state-of-the-art traditional and learning-based image compression algorithms in terms of perceived quality, even at low bitrates. Our method offers a novel perspective for image compression algorithms to improve their quality-aware performance.
2 Related work
@@
End-to-end RS image compression. In RS image compression, many researchers have directed their attention towards network structure and image transformation techniques. For instance, Pan et al. [18] adopted GANs to independently decode the image content and detailed textures, later combining these features to achieve low-bitrate compression of RS images. More recently, Xiang et al. [19] leveraged the discrete wavelet transform to separate image features into high- and low-frequency components. They then designed compression networks to bolster the model’s ability to present these features at different frequencies to recover texture-rich details.
@@
Moreover, in the pursuit of advancing network modules for decoding more texture-rich details of RS image compression, some studies aim to improve the estimated accuracy of the entropy model. Pan et al. [9] proposed a hybrid attention network to improve the prediction accuracy of the entropy model and enhance the representation capabilities of both the encoder and decoder. In [4], a long-range convolutional network was developed as a network component to enhance the decoding performance of RS image compression. Fu et al. [30] considered the local and non-local redundancies presented in RS images and developed a mixed hyperprior network to explore both, thus improving the accuracy of entropy estimation. Besides, Zhang et al. [11] introduced a multi-scale attention module to enhance the network’s feature extraction capability, and employed global priors and anchored-stripe attention to improve the entropy model.
@@
In summary, the above RS image compression algorithms mainly focus on designing complex networks for compact latent variable representation and more accurate entropy estimation to achieve texture-rich decoding results. Different from the above methods, we innovatively project the compression distortion loss of decoded images from an existing image compression algorithm into the Gaussian-distributed latent variables in the forward network, while combining the Gaussian distribution with the decoded images in the inverse network to obtain texture-rich enhanced images.
@@
Invertible neural network (INN). INN offers several advantages, including bijective mapping, efficient mapping access, and computable Jacobi matrices, making them promising for various machine learning tasks [31, 32, 33, 34, 35]. For instance, Zhao et al. [25] employed INN to generate invertible grayscale images by separating color information from color images and encoding it into Gaussian-distributed latent variables. This approach ensures that the color information lost during grayscale generation remains independent of the input image. Similarly, Xiao et al. [36] developed invertible models to generate valid degraded images while transforming the distribution of lost contents to a fixed distribution of a latent variable during forward degradation. Restoration was then made tractable by applying the inverse transformation on the generated degraded image along with a randomly drawn latent variable. Liu et al. [33] considered image degradation and super-resolution as a pair of inverse tasks and replaced the generators in GAN with INN for unpaired image super-resolution. In addition, INN technology has been leveraged to improve the transformation between the image space and the latent feature space, thereby mitigating information loss [23]. For instance, Xie et al. [23] proposed an enhanced INN-based encoding network for better image compression. Gao et al. [26] integrated INN into the development of an invertible image generation module to prevent information loss and developed a competitive low-bitrate image compression algorithm.
3 The proposed INN-RSIC
@@
RS images typically contain more detailed information compared to natural images, posing a greater challenge in obtaining texture-rich decoded RS images, especially at low bitrates [17]. Therefore, our INN-RSIC aims to utilize INN to encode the lost compression distortion of an existing image compression algorithm into a set of latent variables $\mathbf{z}$, following a pre-defined distribution, such as a Gaussian distribution. As a result, the distribution of latent variables becomes independent of the distribution of the input image. In the enhanced reconstruction stage of our framework (i.e., the inverse mapping of our INN-RSIC), a new set of randomly sampled latent variables $\vec{\mathbf{z}}$ can effectively represent the lost compression distortion information to some extent. Thus, we can simply feed the re-sampled $\vec{\mathbf{z}}$, along with the decoded images, to the INN-RSIC of the inverse network for enhanced images.
@@
Specifically, the framework of our INN-RSIC is depicted in Figure 2, consisting of a compressor and an enhancer. Given an input image $\mathbf{x}$, we utilize the competitive image compression algorithm ELIC [15] as the compressor to generate the decoded image $\tilde{\mathbf{x}}$, which serves as a label to constrain the generated synthetic image $\mathbf{s}$ for training the enhancer. The enhancer mainly comprises two streams: the forward and inverse networks, denoted as $f_{\mathbf{\Theta}}(\cdot)$ and $f_{\mathbf{\Theta}}^{-1}(\cdot)$, respectively, where $\mathbf{\Theta}$ represents the network parameters.
Figure 2: Illustration of our INN-RSIC. Compressor denotes the image compression algorithm [15].
3.1 INN architecture
3.1.1 Invertible block
The invertible block layer is a crucial component of invertible architectures, acting as a bridge to adapt between two different distributions via its trainable parameters. The structure of an invertible block is illustrated in Figure 2 (b). The inverse block operates on an input feature map $\mathbf{v}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ denote height, width, and the number of channels, respectively. In each invertible block, $\mathbf{v}$ is splitted along the channel dimension into two parts $[\mathbf{v}_{1},\mathbf{v}_{2}]$, where $\mathbf{v}_{1}\in\mathbb{R}^{H\times W\times C_{1}}$ and $\mathbf{v}_{2}\in\mathbb{R}^{H\times W\times C_{2}}$, and $C_{1}+C_{2}=C$.  To obtain the input features of the $(l+1)$-th invertible block, inverse transformations are applied to both segments using learnable scale and shift parameters by
$\mathbf{v}_{1}^{l+1}=\mathbf{v}_{1}^{l}+\phi(\mathbf{v}_{2}^{l});~{}~{}~{}~{}\mathbf{v}_{2}^{l+1}=\mathbf{v}_{2}^{l}\odot\exp(\rho(\mathbf{v}_{1}^{l+1}))+\eta(\mathbf{v}_{1}^{l+1}),$ (1)
where $\odot$ refers to Hadamard product. The $\varphi(\cdot)$, $\rho(\cdot)$, and $\eta(\cdot)$ are realized by the dense block [37]. Thus, the inverse transformation can be derived from Eq. (1) by
$\mathbf{v}_{2}^{l}=(\mathbf{v}_{2}^{l+1}-\eta(\mathbf{v}_{1}^{l+1}))\oslash\exp(\rho(\mathbf{v}_{1}^{l+1}));~{}~{}~{}~{}\mathbf{v}_{1}^{l}=\mathbf{v}_{1}^{l+1}-\phi(\mathbf{v}_{2}^{l}),$ (2)

### v2

@@ Abstract
Despite the impressive perceptual or fidelity performance of existing image compression algorithms, they typically struggle to strike a balance between perceptual quality and high image fidelity. To address this challenge, we propose a novel invertible neural network-based remote sensing image compression (INN-RSIC) method. Our approach captures the compression distortion of an existing image compression algorithm and encodes it as a set of Gaussian-distributed latent variables via an invertible neural network (INN). This ensures that the compression distortion in the decoded image remains independent of the ground truth. By leveraging the inverse mapping of the INN, we can input the decoded image along with a set of randomly resampled Gaussian-distributed variables into the inverse network, effectively generating enhanced images with improved perceptual quality. To accurately learn and represent compression distortion, we utilize channel expansion, Haar transformation, and invertible blocks in constructing the INN. Additionally, we introduce a quantization module (QM) to mitigate the impact of format conversion, thereby enhancing the framework’s generalization capabilities and improving the perceptual quality of the enhanced images. Extensive experiments demonstrate that the proposed INN-RSIC achieves a superior balance of perceptual quality and high image fidelity compared to existing image compression algorithms. Moreover, the developed INN-based enhancer operates as a plug-and-play (PnP) method, allowing seamless integration with existing image compression algorithms to improve their decoding perceptual quality. Our code and pre-trained models are available at xxxx.
Index Terms:
Remote sensing image compression, invertible neural network, Haar transformation, image enhancement.
I Introduction
@@
Remote sensing (RS) images have been widely applied in various fields, including urban planning [1], biosphere monitoring [2], and land resource management [3]. However, RS images are typically characterized by high content, high resolution, and large size [4, 5]. Furthermore, with the advancement of sensor technology and the enhancement of the image acquisition capability of satellite and airborne equipment, the necessity of storing or transmitting RS images is also growing in importance. To cope with these challenges, high-resolution RS images are usually compressed before they are stored and transmitted to the ground. In this case, it is critical to achieve low-bitrate compression while obtaining the decoded images with high perceptual quality.
@@
Traditional image compression algorithms, including JPEG2000 [6], BPG [7], and VVC [8], have been instrumental in facilitating the storage and transmission of image data. However, image compression with these algorithms inevitably suffered from undesired blocking, ringing artifacts, and blurring [9, 10], which may highly influence the perception quality.
@@
With the development of deep learning techniques, learning-based image compression methods have made significant progress in obtaining high rate-distortion (RD) performance [11, 9, 12, 13, 14, 15, 16, 17]. Notably, in natural image compression, Ballé et al. [18] proposed to view additional side information as a hyper-prior entropy model to estimate a zero-mean Gaussian distribution, laying the foundation for subsequent improvements in entropy modeling. Cheng et al. [10] further introduced discretized Gaussian mixture likelihoods to realize the distribution estimation of the entropy model, resulting in impressive decoding performance. Meanwhile, He et al. [19] developed a spatial-channel contextual adaptive model, enhancing compression performance without sacrificing computational speed. Liu et al. [16] leveraged the local modeling ability of convolutional neural networks (CNNs) and the non-local modeling strengths of Transformers to develop the encoder and decoder networks. A channel-squeezing-based entropy model was further proposed to enhance RD performance. Jiang et al. [17] proposed capturing the channel-wise, local spatial, and global spatial correlations present in latent representations to develop a comprehensive entropy model, which is employed to design the competitive image compression method MLIC+. Despite the success of learning-based compression algorithms in natural scenes, RS images present unique challenges due to their rich texture, weak correlation, and low redundancy compared to natural images [20, 21]. Preserving visually pleasing decoded images while achieving high compression bitrates remains a significant challenge in RS image compression.
@@
To tackle the challenge, Zhang et al. [11] introduced a multi-scale attention module to enhance the network’s feature extraction capability, and developed an improved entropy model by using global priors and anchored-stripe attention. Pan et al. [22] resorted to generative adversarial networks (GANs) to independently reconstruct the image content and detailed textures, subsequently fusing these features to achieve low-bitrate RS image compression. Additionally, Xiang et al. [23] utilized discrete wavelet transform to separate image features into high- and low-frequency components, and designed compression networks to enhance the model’s representation ability of both types of features.
@@
In summary, those above deep learning-based image compression methods typically achieve image compression through CNN blocks [10, 19, 18, 13, 11, 23] or Transformer blocks [24, 25, 26, 27, 28], followed by optimization using distance measurements like L1, L2, or adversarial loss [29]. Despite their impressive performance, these methods usually struggle to generate satisfactory perception quality while maintaining high image fidelity. Recently, invertible neural networks (INN) have emerged as a new paradigm for image generation, demonstrating remarkable performance across various applications [30, 31, 32, 33]. Building on this success, several approaches have integrated INN into image compression to capture richer texture details [27, 31]. Fig. 1 illustrates the overview of existing INN-based compression methods. For instance, in [27], an enhanced INN-based encoding network was devised to improve the decoding performance of natural image compression. In [31], the authors utilized INN to develop an invertible image generation module, aiming to prevent information loss and propose a competitive low-bit-rate compression algorithm. Unlike these methods, the proposed approach focuses on leveraging the invertible capabilities of INN to recover more texture information of images decoded by an existing compression algorithm, without requiring additional bitstreams and model retraining. It is expected to sample some texture details from Gaussian distributed variables conditioned on the decoded image. Fig. 2 illustrates the perception comparison between the decoded images and the enhanced images. One can see that the proposed INN-RSIC exhibits a high capability in recovering texture-rich details without additional bitrates.
Figure 1: Overview of existing INN-based image compression methods. For simplicity, entropy models are omitted. It can be seen that our method is a plug-and-play (PnP) approach that does not require retraining of existing compression models. Our method can obtain both decoded images with high PSNR and MS-SSIM output from the original compression algorithm (e.g., ELIC) and enhanced images with high perceptual quality.
Figure 2: Visualization of the decoded images by ELIC [19] and the proposed INN-RSIC.
@@
In this paper, we propose the invertible neural network-based remote sensing image compression (INN-RSIC). Specifically, we utilize an invertible forward network with a conditional generation module (CGM) to encode compression loss information of an existing image compression algorithm into Gaussian-distributed latent variables. As a result, it is expected to sample some of the lost prior information from Gaussian distribution, thereby facilitating the recovery of visually enhanced images during the inverse mapping. Additionally, to effectively learn compression distortion, we adopt channel expansion and Haar transformation [34] to separate high and low frequency and introduce a quantization module (QM) to reduce the impact of the reconstruction quality due to data type conversion in the inference stage.
The primary contributions of this paper can be summarized as follows:
  • $\bullet$
    The proposed INN-RSIC is the first attempt to model image compression distortion of an image compression method using invertible transforms. It serves as a PnP method to obtain highly perceptible decoded images while preserving the performance of the baseline algorithm.
  • $\bullet$
    We develop a novel, effective yet simple architecture with channel expansion, Haar transformation, and invertible blocks. This architecture enables projecting compression distortion into a case-agnostic distribution, so that compression distortion information can be obtained based on samples from a Gaussian distribution.
  • $\bullet$
    CGM is introduced to split and encode the compression distortion information from ground truth into the latent variables conditioned on the synthetic image. This process generates pattern-free synthetic images and enhances the texture details of the reconstructed images during the inverse mapping.
  • $\bullet$
    Extensive experiments indicate that the proposed INN-RSIC achieves a superior balance of perceptual quality and high image fidelity compared to existing state-of-the-art image compression algorithms. Our method offers a novel perspective for image compression algorithms to improve their perception quality.
II Related work
II-A End-to-End RS Image Compression
@@
In RS image compression, many researchers have directed their attention towards network structure and image transformation techniques. For instance, Pan et al. [22] adopted GANs to independently decode the image content and detailed textures, later combining these features to achieve low-bitrate compression of RS images. More recently, Xiang et al. [23] leveraged the discrete wavelet transform to separate image features into high- and low-frequency components. They then designed compression networks to bolster the model’s ability to present these features at different frequencies to recover texture-rich details.
@@
Moreover, in the pursuit of advancing network modules for decoding more texture-rich details of RS image compression, some studies aim to improve the estimated accuracy of the entropy model. Pan et al. [9] proposed a hybrid attention network to improve the prediction accuracy of the entropy model and enhance the representation capabilities of both the encoder and decoder. In [4], a long-range convolutional network was developed as a network model to improve the decoding performance of RS image compression. Fu et al. [35] considered the local and non-local redundancies presented in RS images and developed a mixed hyperprior network to explore both, thus improving the accuracy of entropy estimation.
@@
In summary, the above RS image compression algorithms mainly focus on designing complex networks for compact latent variable representation and more accurate entropy estimation to achieve texture-rich decoding results. Different from the above methods, we innovatively project the compression distortion loss of decoded images from an image compression method into Gaussian-distributed latent variables in the forward network, while combining the Gaussian distribution with the decoded images in the inverse network to obtain texture-rich enhanced images.
II-B Invertible Neural Network (INN)
@@
INN offers several advantages, including bijective mapping, efficient mapping access, and computable Jacobi matrices, making them promising for various machine learning tasks [36, 37, 38, 39, 40, 41, 42]. For instance, Zhao et al. [30] employed INN to generate invertible grayscale images by separating color information from color images and encoding it into Gaussian-distributed latent variables. This approach ensures that the color information lost during grayscale generation remains independent of the input image. Similarly, Xiao et al. [43] developed invertible models to generate valid degraded images while transforming the distribution of lost contents to a fixed distribution of a latent variable during forward degradation. The restoration was then made tractable by applying the inverse transformation on the generated degraded image along with a randomly drawn latent variable. Liu et al. [39] considered image degradation and super-resolution as a pair of inverse tasks and replaced the generators in GAN with INN for unpaired image super-resolution. In addition, INN technology has been leveraged to improve the transformation between the image space and the latent feature space, thereby mitigating information loss [27]. For instance, Xie et al. [27] proposed an enhanced INN-based encoding network for better image compression. Gao et al. [31] integrated INN into the development of an invertible image generation module to prevent information loss and developed a competitive low-bitrate image compression algorithm.
III The Proposed INN-RSIC
@@
Compared to natural images, RS images usually contain richer textures, which makes acquiring RS images at low bit rates more challenging [21]. Therefore, the proposed INN-RSIC aims to utilize INN to encode the lost compression distortion of an existing image compression algorithm into a set of latent variables $\mathbf{z}$, following a pre-defined distribution, such as a Gaussian distribution. As a result, the distribution of latent variables becomes independent of the distribution of the input image. In the enhanced reconstruction stage of our framework (i.e., the inverse mapping of the proposed INN-RSIC), a new set of randomly sampled latent variables $\vec{\mathbf{z}}$ can effectively represent the lost compression distortion information to some extent. Thus, we can simply feed the re-sampled $\vec{\mathbf{z}}$, along with the decoded images, to the INN-RSIC of the inverse network for enhanced images.
@@
Specifically, the architecture of the proposed INN-RSIC is depicted in Fig. 3, consisting of a compressor and an enhancer. Given an input image $\mathbf{x}$, we utilize the competitive image compression algorithm ELIC [19] as the compressor to generate the decoded image $\tilde{\mathbf{x}}$, which serves as a label to constrain the generated synthetic image $\mathbf{s}$ for training the enhancer. The enhancer mainly comprises two streams: the forward and inverse networks, denoted as $f_{\mathbf{\Theta}}(\cdot)$ and $f_{\mathbf{\Theta}}^{-1}(\cdot)$, respectively, where $\mathbf{\Theta}$ represents the network parameters.
Figure 3: The architecture of the proposed INN-RSIC. Compressor refers to the competitive image compression algorithm [19].
III-A INN Architecture
III-A1 Invertible Block
The invertible block layer is a crucial component in invertible architectures, acting as a bridge to adapt between two different distributions via its trainable parameters. Fig. 3(b) depicts the structure of an invertible block. An inverse block operates on an input feature map $\mathbf{v}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ denote height, width, and the number of channels, respectively. In each invertible block, $\mathbf{v}$ is splitted along the channel dimension into two parts $[\mathbf{v}_{1},\mathbf{v}_{2}]$, where $\mathbf{v}_{1}\in\mathbb{R}^{H\times W\times C_{1}}$ and $\mathbf{v}_{2}\in\mathbb{R}^{H\times W\times C_{2}}$, and $C_{1}+C_{2}=C$.  To obtain the input features of the $(l+1)$-th invertible block, inverse transformations are applied to both segments using learnable scale and shift parameters by
$\mathbf{v}_{1}^{l+1}$ $=\mathbf{v}_{1}^{l}+\phi(\mathbf{v}_{2}^{l}),$ (1)
$\mathbf{v}_{2}^{l+1}$ $=\mathbf{v}_{2}^{l}\odot\exp(\rho(\mathbf{v}_{1}^{l+1}))+\eta(\mathbf{v}_{1}^{l+1}),$
where $\odot$ refers to Hadamard product. The $\varphi(\cdot)$, $\rho(\cdot)$, and $\eta(\cdot)$ are achieved by the dense blocks [44]. Therefore, the inverse transformation can be derived from Eq. (1) by
$\mathbf{v}_{1}^{l}$ $=\mathbf{v}_{1}^{l+1}-\phi(\mathbf{v}_{2}^{l}),$ (2)
$\mathbf{v}_{2}^{l}$ $=(\mathbf{v}_{2}^{l+1}-\eta(\mathbf{v}_{1}^{l+1}))\oslash\exp(\rho(\mathbf{v}_{1}^{l+1})),$